# NID-SLAM: Neural Implicit Representation-based RGB-D SLAM in dynamic environments

### v1

@@ Abstract
Index Terms—  Neural implicit representation, semantic segmentation, visual SLAM
1 Introduction
@@
Recent methods [2] have introduced neural implicit representation into SLAM. The most notable example, neural radiance fields (NeRF) [3], encodes scene color and volume density into the weights of a neural network, directly learning high-frequency scene details from the data, significantly enhancing mapping smoothness and coherence. Combined with rendering methods based on volumetric representation, NeRF, through training, can re-synthesize input images and generalize to adjacent, unseen viewpoints. However, most neural SLAM algorithms are based on the assumption of a static environment, with some capable of handling small dynamic objects in synthetic scenes [4]. In real-world dynamic scenarios, these algorithms could experience a significant performance-degradation in terms of dense reconstruction and camera tracking accuracy. This could be largely caused by the existence of dynamic objects in the scene that induce incorrect data associations, thereby severely disrupting the pose estimation during the tracking process. Additionally, information about dynamic objects is often incorporated into the map, hindering its long-term applicability.
Fig. 1: The 3D reconstruction results of NID-SLAM on our self-captured large dynamic scene.
@@ 2 RELATED WORK
Neural implicit represetations [3, 10, 11], which encode the 3D geometry and appearance of scenes into the weights of neural networks, have gained popularity due to their compact and precise results. When applied to visual SLAM, this technique enables the construction of environmental maps and the estimation of camera poses. Previous methods [2, 4, 12, 13, 14] have demonstrated the feasibility of using neural networks to model the color and geometric information of static scenes; however, they have not fully exploited the potential of neural implicit representations in dynamic environments. Instead of the approach used in [4], which resolves small dynamic objects in synthetic scenes by discarding outliers in losses, we employ a comprehensive removal method to address large dynamic objects in real-world scenes.
Fig. 2: System overview. 1) Dynamic object removal: by employing semantic segmentation and mask revision, dynamic objects within RGB-D images are precisely eliminated, followed by a thorough restoration of the occluded backgrounds. 2) Tracking: camera poses $\{\mathbf{R},\mathbf{t}\}$ are optimized by minimizing losses. 3) Mapping: a mask-guided strategy is employed to select keyframes for the optimization of feature grids scene representation. 4) Scene representation: efficient rendering of predicted color and depth values is achieved through surface-focused point sampling.
3 Methods
@@
3.1 Dynamic object removal
@@
3.2 Mask-guided keyframe selection
@@
3.3 Scene representation & image rendering
@@ 3.4 Mapping and tracking
8:     Predict color $c_{\mathbf{x}}$ for sampled points $\mathbf{x}$ by Eq. (2)
9:     Render $R_{i}$ by Eq. (3)
@@
4 Experiments
Table 1: Camera tracking results on TUM RGB-D. ATE RMSE $[m](\downarrow)$ is used as evaluation metric. \ represents corresponding data is not mentioned in the respective literature.
@@ 4 Experiments
Table 2: Translational RPE RMSE results $[m](\downarrow)$ on TUM.
@@
Table 3: Rotational RPE RMSE results $[m](\downarrow)$ on TUM.
@@
${}^{\star}$ w mask Refer to caption
@@
Fig. 3: Reconstruction results on TUM RGB-D. The red box highlights areas with dynamic object.
4.1 Experimental setup
@@
Metrics. We evaluate camera tracking using Absolute Trajectory Error (ATE) and Relative Pose Error (RPE). For evaluation of the reconstruction quality, we adopt Depth L1 (cm), Accuracy (cm), Completion (cm), and Completion ratio (%) with a threshold of 5cm. Unlike NICE-SLAM [4], we present results without extra mesh culling.
Implementation details. We run NID-SLAM on a desktop PC with a 3.20GHz Intel Core i9-12900KF CPU and NVIDIA RTX 3090 GPU. Note that ${}^{\star}$ denotes the iMAP reimplementation released by the NICE-SLAM [4] authors.
4.2 Evaluation of mapping and tracking
Evaluation on TUM RGB-D [16]. To evaluate camera tracking performance, we use the dynamic TUM RGB-D dataset. For fair comparison, we provided ${}^{\star}$ and NICE-SLAM with the same dynamic-object-free input as our method. As shown in Table 1, our approach achieves the best ATE results among neural SLAM methods and surpasses ORB-SLAM2 in several scenarios. Table 2 and Table 3 present the result of translational RPE and rotational RPE, our method significantly outperforms other neural SLAM approaches. However, due to the inherent gap in tracking performance between neural SLAM methods and classic SLAM approaches, NID-SLAM is not yet able to outperform Dyna-SLAM. Nevertheless, our method significantly narrows the performance gap in tracking between these two categories of methods in dynamic environments. Figure 3 qualitatively presents the reconstruction results on TUM RGB-D. Notably, our approach achieves accurate removal of dynamic objects, significantly improving the reconstruction quality and completeness of dynamic scenes.
Evaluation on Replica [17]. We also evaluate the reconstruction quality on Replica while using the same rendered RGB-D sequence provided by the authors of NICE-SLAM. With the keyframe selection method, our method ensures high-quality reconstruction across both central and edge regions. In Table 4, we present a quantitative comparison of the reconstruction and rendering performance of our method and the baselines. The results demonstrate that our approach outperforms the baselines for both 2D and 3D metrics. Qualitatively, we can see from Fig. 3 that the mesh obtained by our method features notably smoother floors and more details on small objects, such as table lamps and ornaments.
@@ 4.2 Evaluation of mapping and tracking
Table 4: Reconstruction results in the Replica dataset (average over 8 scenes).
${}^{\star}$ DI-Fusion [18] NICE-SLAM Ours
@@
Fig. 4: Reconstruction results on Replica dataset. The red box highlights the improved areas.
Table 5: Ablation study of our design choices on the TUM sequence fr3/w/half and fr3/w/xyz.
@@
4.3 Ablation study
Depth revision. We demonstrate the effectiveness of depth revision on TUM [16]. As shown in Table 5, failing to revise depth information and further refining semantic masks not only significantly reduces tracking accuracy but also leads to less robustness.
Keyframe selection. We test our method using overlap-based strategy (as shown in the second row of Table 5), where they simply perform equidistant selection. In contrast, our specially designed strategy for dynamic scenes not only enhances tracking robustness but also enriches the keyframe set with extensive scene information, playing a vital role in accurate tracking.
Sampling. In Table 5, we also illustrate the effectiveness of our pixel sampling strategy. Results indicate that performing surface-based sampling and filtering out irrelevant points for rendering enhances the quality of sampled points, thereby improving the accuracy of camera tracking.
5 Conclusion

### v2

@@ Abstract
Index Terms:
neural implicit representation, semantic segmentation, visual SLAM
Figure 1: The 3D reconstruction results of NID-SLAM on our self-captured large dynamic scene.
I Introduction
@@
Recent methods [2] have introduced neural implicit representation into SLAM. The most notable example, neural radiance fields (NeRF) [3], encodes scene color and volume density into the weights of a neural network, directly learning high-frequency scene details from the data, significantly enhancing mapping smoothness and coherence. Combined with rendering methods based on volumetric representation [4], NeRF, through training, can re-synthesize input images and generalize to adjacent, unseen viewpoints. However, most neural SLAM algorithms are based on the assumption of a static environment, with some capable of handling small dynamic objects in synthetic scenes [5]. In real-world dynamic scenarios, these algorithms could experience a significant performance-degradation in terms of dense reconstruction and camera tracking accuracy. This could be largely caused by the existence of dynamic objects in the scene that induce incorrect data associations, thereby severely disrupting the pose estimation during the tracking process. Additionally, information about dynamic objects is often incorporated into the map, hindering its long-term applicability.
@@ II RELATED WORK
Neural implicit represetations [3, 11, 12], which encode the 3D geometry and appearance of scenes into the weights of neural networks, have gained popularity due to their compact and precise results. When applied to visual SLAM, this technique enables the construction of environmental maps and the estimation of camera poses. Previous methods [2, 5, 13, 14, 15, 16] have demonstrated the feasibility of using neural networks to model the color and geometric information of static scenes; however, they have not fully exploited the potential of neural implicit representations in dynamic environments. Instead of the approach used in [5], which resolves small dynamic objects in synthetic scenes by discarding outliers in losses, we employ a comprehensive removal method to address large dynamic objects in real-world scenes.
Figure 2: System overview. 1) Dynamic object removal: by employing semantic segmentation and mask revision, dynamic objects within RGB-D images are precisely eliminated, followed by a thorough restoration of the occluded backgrounds. 2) Tracking: camera poses $\{\mathbf{R},\mathbf{t}\}$ are optimized by minimizing losses. 3) Mapping: a mask-guided strategy is employed to select keyframes for the optimization of feature grids scene representation. 4) Scene representation: efficient rendering of predicted color and depth values is achieved through surface-focused point sampling.
III Methods
@@
III-A Dynamic object removal
@@
III-B Mask-guided keyframe selection
@@
III-C Scene representation & image rendering
@@ III-D Mapping and tracking
8:     Predict color $c_{\mathbf{x}}$ for sampled points $\mathbf{x}$ by (2)
9:     Render $R_{i}$ by (3)
@@
IV Experiments
TABLE I: Camera tracking results on TUM RGB-D. ATE RMSE $[m](\downarrow)$ is used as evaluation metric. \ represents corresponding data is not mentioned in the respective literature.
@@ IV Experiments
TABLE II: Translational RPE RMSE results $[m](\downarrow)$ on TUM.
@@
TABLE III: Rotational RPE RMSE results $[m](\downarrow)$ on TUM.
@@
iMAP w mask Refer to caption
@@
Figure 3: Reconstruction results on TUM RGB-D. The red box highlights areas with dynamic object.
IV-A Experimental setup
@@
Metrics. We evaluate camera tracking using Absolute Trajectory Error (ATE) and Relative Pose Error (RPE). For evaluation of the reconstruction quality, we adopt Depth L1 (cm), Accuracy (cm), Completion (cm), and Completion ratio (%) with a threshold of 5cm.
Implementation details. We run NID-SLAM on a desktop PC with a 3.20GHz Intel Core i9-12900KF CPU and NVIDIA RTX 3090 GPU. In all our experiments, we set the number of sampling points on a ray to $M_{\text{strat}}=32$ and $M_{\text{surf}}=16$; we select $K=5$ keyframes and sample $N=1000$ pixels for optimization, along with $N_{t}=200$ pixels for tracking. Note that iMAP denotes the iMAP reimplementation released by the NICE-SLAM [5] authors.
IV-B Evaluation of mapping and tracking
Evaluation on TUM RGB-D [18]. To evaluate camera tracking performance, we use the dynamic TUM RGB-D dataset. For fair comparison, we provided iMAP and NICE-SLAM with the same dynamic-object-free input as our method. As shown in Table I, our approach achieves the best ATE results among neural SLAM methods and surpasses ORB-SLAM2 in several scenarios. Table II and Table III present the result of translational RPE and rotational RPE, our method significantly outperforms other neural SLAM approaches. However, due to the inherent gap in tracking performance between neural SLAM methods and classic SLAM approaches, NID-SLAM is not yet able to outperform Dyna-SLAM. Nevertheless, our method significantly narrows the performance gap in tracking between these two categories of methods in dynamic environments. Figure 3 qualitatively presents the reconstruction results on TUM RGB-D. Notably, our approach achieves accurate removal of dynamic objects, significantly improving the reconstruction quality and completeness of dynamic scenes.
Evaluation on Replica [19]. We also evaluate the reconstruction quality on Replica while using the same rendered RGB-D sequence provided by the authors of NICE-SLAM. With the keyframe selection method, our method ensures high-quality reconstruction across both central and edge regions. In Table IV, we present a quantitative comparison of the reconstruction and rendering performance of our method and the baselines. The results demonstrate that our approach outperforms the baselines for both 2D and 3D metrics. Qualitatively, we can see from Fig. 3 that the mesh obtained by our method features notably smoother floors and more details on small objects, such as table lamps and ornaments.
@@ IV-B Evaluation of mapping and tracking
TABLE IV: Reconstruction results in the Replica dataset (average over 8 scenes).
iMAP DI-Fusion [20] NICE-SLAM Ours
@@
Figure 4: Reconstruction results on Replica dataset. The red box highlights the improved areas.
TABLE V: Ablation study of our design choices on the TUM sequence fr3/w/half and fr3/w/xyz.
@@
IV-C Ablation study
Depth revision. We demonstrate the effectiveness of depth revision on TUM [18]. As shown in Table V, failing to revise depth information and further refining semantic masks not only significantly reduces tracking accuracy but also leads to less robustness.
Keyframe selection. We test our method using overlap-based strategy (as shown in the second row of Table V), where they simply perform equidistant selection. In contrast, our specially designed strategy for dynamic scenes not only enhances tracking robustness but also enriches the keyframe set with extensive scene information, playing a vital role in accurate tracking.
Sampling. In Table V, we also illustrate the effectiveness of our pixel sampling strategy. Results indicate that performing surface-based sampling and filtering out irrelevant points for rendering enhances the quality of sampled points, thereby improving the accuracy of camera tracking.
V Conclusion
@@
Acknowledgment
This work was partially supported by the Key R&D Program of Zhejiang Province (No. 2023C01181)